Consider an economy having following values of Consumption, Investment, Government

 

Spending, and Taxes, 

 

C = 300 + 0.7Yd         I = 200 – 200i G = 600          T = 100    

 

Derive the aggregate demand equation. Also, calculate the output in an economy


1
Expert's answer
2021-06-02T11:58:23-0400

Solution:

Aggregate demand equation: AE = Y

Aggregate demand equation = C + I + G

T = 100

C = 300 + 0.7Yd = 300 + 0.7(Y – T) = 300 + 0.7(Y – 100) = 300 +0.7Y – 70 = 300 – 70 + 0.7Y

C = 230 + 0.7Y

I = 200 – 200i

G = 600

Aggregate demand equation = 230 + 0.7Y + 200 – 200i + 600

 

Output (Y) = C + I + G

Y = 230 + 0.7Y + 200 – 200i + 600

Y – 0.7Y = 230 + 200 + 600 – 200i

0.3Y = 1,030 – 200i

Divide both sides by 0.3:

Y = 3,433.33 – 666.67i

 

Output (Y) = 3,433.33 – 666.67i